2.10. Let H be the subgroup generated by two elements a, b of a group G. Prove that if ab = ba, then H is an abelian group.


1
Expert's answer
2022-06-22T12:32:00-0400

The elemtnts of "H" are of the form: "a^{i_1}b^{i_2}a^{i_3}\cdots a^{i_k-1}b^{i_k}" where "i_1,\dots, i_k\in\mathbb{Z}" for some "k".


So let "x,y\in H". Then we can write "x=a^{i_1}b^{i_2}a^{i_3}\cdots a^{i_k-1}b^{i_k}" and "y=a^{j_1}b^{j_2}a^{j_3}\cdots a^{j_k-1}b^{j_k}". Then "xy=(a^{i_1}\cdots b^{i_k})(a^{j_1}\cdots b^{j_k})" since "ab=ba" we can interchange each term in this multiplication, and obtain: "xy=(a^{j_1}\cdots b^{j_k})(a^{i_1}\cdots b^{i_k})=yx".

This implies "H" is abelian.